The following data gives the number of defectives in 5 independent samples from a production process. The samples are of varying size. Draw appropriate chart and comment on the process

Sample number     Sample size     Number of defectives

1                              2000                   400

2                              1000                   150

3                              1000                   120

4                               600                     80

5                               400                     50
Answer #1

For our example we use np or do chart For this chart it is not necessary to calculate the problem of proportion of defectives[LCLd = 122.665] Cld = npe 5000 x 0:032 = 160 [cLd=160 UCLI = np + 3 vno a = ( 5600 X 0.032) + 3 J5000X 0.03 2x0.968 = 160+ 3d-chart is scale on axis = 1cm= 1 on-y-axis = 50 detectives 250 Defectives 2.00 150 JUCL → CL LCL TO 0 1 2 3 4 5 sample numbe